OpenStudy (anonymous):

What is the equation for a circle Center (-2, -5) radius =1

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

thanks but not quite what i was looking for

OpenStudy (anonymous):

(h,k)

OpenStudy (anonymous):

your h is -2, k is -5

OpenStudy (anonymous):

so you plug it in then square your radius

OpenStudy (anonymous):

\[(x+2)^2+(y+5)^2=1^2\]

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

HOW BOUGHT this one with center (0,8) radius =8

OpenStudy (anonymous):

well i showed you how to do it, wanna give it a shot?

OpenStudy (anonymous):

its really just plugging in

OpenStudy (anonymous):

ok so ( x + 0)2 + ( y+8)2 = 8 2

OpenStudy (anonymous):

the only reason why i changed it to positive the first time was because the numbers were negative, so for the most part, you're correct!

OpenStudy (anonymous):

\[x^2+(y-8^2)=8^2\]

OpenStudy (anonymous):

Center ( 7, -3) radius =7

OpenStudy (anonymous):

use the same formula